Angus invested 18,000, part at 3% and part at 5%. If the total interest at the end of the year is 660, how much did he invest at each rate?
Answer by checkley71(8403)

.03x+.05(18000-x)=660 or .03x+900-.05x=660 or -.02x=-240 or x=240/.02 or x=12000 invested @ 3% then we have 18000-12000=6000 invested @ 5%
proof .03*12000+.05*6000=660 or 360+300=660 or 660=660
YOU ARE CORRECT 12,000*.03 REPRESENTS $12,000 INVESTED @ 3%
AND 6,000*.05 REPRESENTS $6,000 INVESTED @5%.